# Traffic Modeling for Network Security and Privacy: Challenges Ahead

### v2

@@ Abstract
Network traffic analysis using machine learning (ML) made significant progress over the past decades. Traffic analysis addresses various challenging problems in network security, ranging from detection of anomalies and attacks to countering of censorship. ML models are also developed to expose user privacy risks as demonstrated by the research works on fingerprinting of user-visiting websites, IoT devices, and different applications, even when payloads are encrypted.
Despite these advancements, significant challenges remain in the domain of network traffic analysis to effectively secure our networks from evolving threats and attacks. After briefly reviewing the relevant tasks and recent ML models for traffic analysis, we discuss the challenges that lie ahead.
@@ 1 Introduction
The sophistication of networks attacks has increased over time due to the rapid proliferation of new technologies, applications, network protocols and devices. Meanwhile, network traffic keeps increasing, both in volume and speed, further complicating security challenges. To effectively secure the networks, traffic analysis has long been recognized as a critical component. As payloads in network packets are hardly available for signature-based detection with the high adoption of TLS by different applications (Google-transparency-mail; Google-transparency-HTTP), statistical models are becoming inevitable to analyze network traffic. Decades of research efforts have led to the development of various statistical and machine learning (ML) models of different capabilities to tackle multiple network security and privacy (NetS&P) tasks.
Deep learning (DL) models, in particular, are valuable for modeling network traffic for NetS&P tasks for two key reasons, or rather, to address two significant challenges. First, network data is enormous; for example, a 1 Gbps link generates hundreds of gigabytes of data per day under high utilization, even when considering only packet headers (excluding the payload). And network bandwidth has increased steadily, with even consumer (broadband) bandwidth now reaching 1 Gbps. Meanwhile, enterprises and telcos/ISPs utilize bandwidth capacities that are orders of magnitude higher. Second, the number of features that can be extracted from network traces can potentially reach several hundreds or even thousands in count. For instance, when modeling a user browsing session as a data point for training/inference, a small session can easily consist of 200 packets, with each packet containing 10 attributes (e.g., packet size, inter-arrival time, and so on). Consequently, representing this relatively short session requires 2,000 features. As we know, DL models are able to process and learn useful patterns from huge datasets with large feature space.
In light of the current state-of-the-art of machine learning in network traffic modeling, this works aims to identify the next set of challenges that lie ahead in securing networks and protecting user privacy.
@@ 2 NetS&P: Tasks and Models
We discuss important NetS&P tasks and model-based solutions in the literature, without aiming for exhaustiveness.
@@ 2.1 Anomaly detection
ML models: The primary challenge for this task stems from the diverse types of network anomalies, including those that are previously unknown. For instance, while the recent encrypted DNS protocols like DoH provide better privacy for users, they also offer malware developers new ways to hide their communications (survey-encrypted-DNS-mw-2022), say, for exfiltrating sensitive information (real-time-DNS-exfiltration-NDSS-2024). These are relatively new anomalies unknown before DoH was deployed. Existing research in literature tackles this challenge of anomaly detection by developing unsupervised and semi-supervised models. Early research works explored PCA (Principal Component Analysis), statistical hypothesis testing and regression models to detect anomalous patterns in network traffic (PCA-Diot-2004; brauckhoff2009applying; NADA-2018; evidence-gathering-2017). With the emergence of generative deep learning models for unsupervised problem settings (AD-robust-AE-2017; adversarially-learned-AD-ICDM-2018), we witnessed the development of reconstruction-based unsupervised network anomaly detection solutions in the last decade (KitSUne-2018-NDSS; GEE-2019). For example, GEE (GEE-2019) trains a variational autoencoder (VAE) on noisy benign network traffic sessions to learn its corresponding representation in the latent space, which it subsequently uses to detect deviating behaviors during the inference stage.
@@ 2.2 Attack classification
ML models: While the problem might appear to be a multi-class classification problem, the data available for different attacks vary. DDoS requires simple features that capture the traffic rate, whereas detecting the presence of malware that uses DGA (domain generation algorithm) requires analysis of (plain-text) DNS payload. Therefore, solutions in this space typically train binary classifiers to detect specific attacks. For example, several ML algorithms for DDoS detection are evaluated considering both accuracy and inference latency in (DDoS-eval-latency-ICNP-2024). Similarly, prior works have proposed methods for classifying DGA domains (DGA-Rossow-2024), identifying bot traffic to servers using GAN-based data augmentation (botnet-S&P-2020), and detecting encrypted malware communications over TLS (anderson2017machine) as well as Tor (Tor-malware-CCS-2022).
@@ 2.3 IoT device identification
ML models: IoT device identification has received considerable attention over the past decade, resulting in the proposal of various supervised and semi-supervised approaches for classifying known and unknown devices, respectively. They include several works that are based on conventional ML as well as DL algorithms (2017-profileiot; 2019-DIoT; DEFT-2019; IoT-biLSTM-2020; WU-ZEST). A recent work (usenix-2020-youarewhatyou) demonstrates that devices connected to a WiFi network can be identified by training a DL model on the broadcast/multicast traffic observed.
@@ 2.4 Website fingerprinting (WFP) attack
ML models: Given that a browsing session consists of hundreds or even thousands of packets, this topic has witnessed significant development of deep learning-based attack strategies. In other words, the ability of deep learning models to learn from a large feature space represents a substantial advantage for this task. The literature has interesting WFP attack proposals that use various DL architectures ranging from CNN to AE to sequence models such as LSTM and transformer (NDSS-2018-wfp-lstm; pfp-PETS-2019; varCNN-WFP-2019; Usenix-2022-onlinewebsitefingerprinting; webpage_quic; multi-tab-FP-transformer-SP2023; transformer-WFP-CCS-2023; MIL-WFP-USENIX-2024; Levi-2024). There are also strategies proposed for countering them (gong2020FRONTGLUE; smith2022QCSD; GAN-WFP-defense-SP-2022; siby2023PET; Usenix-2023-website-fingerprinting). These models are evaluated under two scenarios: i) a closed-world (lab) setting, where a classification model is trained to classify $n$ specific monitored websites (each with samples collected over time) and is tested for accurate classification of the same $n$ websites, where $n$ is typically in a few hundreds; ii) a more realistic open-world setting, where the challenge is to identify $n$ monitored websites under the assumption that the user visits $m~(m\gg n)$ unmonitored websites, where $m$ is in tens of thousands. Since tunneling of traffic hides packet headers, (different from the previous tasks) the features used here are the basic (or raw) ones such as packet-size, inter-arrival time (IAT) (or time lag) of packet and direction of packet (as browsing generates bidirectional traffic).
@@ 2.5 Other tasks and models
We also highlight the application of adversarial ML, which aims to generate adversarial samples for evading systems utilizing ML models. In the context of NetS&P, an adversarial ML approach can be exploited not only to bypass ML-based security systems but also to counter threats related to privacy and censorship. For instance, authors in (shenoi2023ipet) adapt generative adversarial perturbations (poursaeed2018generative) for evading IoT fingerprinting solutions that employ ML models. And in the space of website fingerprinting, a defense based on GAN, specifically, WGAN-div (wu2018wasserstein), to evade WFP attacker (discriminator) is studied in (GAN-WFP-defense-SP-2022).
@@ 3.1 Data challenges
One of the most important questions is, do the current model-based solutions generalize beyond the experiments conducted in controlled lab environments? Addressing this question is challenging due to the lack of high-quality labeled data. Most research on network anomaly detection, attack classification, and device identification relies on openly available labeled datasets. However, since these datasets are generated in lab environments, they often contain artifacts or ‘bad design smells’ (bad-smells-EuroSP-2024) that raise concerns on model over-fitting and biases. In  (bad-smells-EuroSP-2024), the authors analyze seven highly-cited datasets and provide evidence of such artifacts in network intrusion detection research. For instance, in one dataset with eight attack classes, two basic features exhibit minimal overlap between benign and attack traffic. As a result, a simple perturbation of these two features—unrelated to the inherent characteristics of attacks—can evade multiple ML and DL models. This simple approach outperforms a state-of-the-art adversarial attack (adversarial-NIDS-2022) employing evolutionary computation and generative adversarial networks. This finding suggests that evaluations in (adversarial-NIDS-2022) may not provide a “meaningful measure of the attack’s effectiveness” (bad-smells-EuroSP-2024) due to dataset flaws. Similar biases in network datasets have also been highlighted for other NetS&P tasks, e.g., website fingerprinting (data-explainable-WFP-PETS-2023; retrace-Tor-2024; genuineTorTraces-2024).
The issue at hand is not about the existing proposals or their evaluation strategies, but rather the quality of the datasets used for evaluating network security tasks. Going one step further, the fundamental challenge is the difficulty of validating the quality network traffic. The traffic generated by a single day’s communications in an enterprise network can reach hundreds of gigabytes, and when considering this scale over multiple days, across various enterprises and different verticals, the volume becomes staggering. Therefore, collecting and labeling real-life network traffic flows is an obstacle. Furthermore, real-world network traffic contains sensitive user information, and sharing this data poses significant privacy risks and ethical concerns. Model-based attacks can extract hidden patterns as fingerprinting attacks in the previous section reveal; meta-data from packet headers (excluding payloads) is sufficient to infer sensitive information. Consequently, it is understandable that enterprises and telco operators refrain from publicly sharing traffic collected from their networks for research and development purposes.
@@ 3.2 Practical deployment
Real-time per-packet inference in network traffic analysis faces several practical limitations. Enterprise networks operate at tens of Gbps whereas telco networks operate at 100s of Gpbs. Even at 10 Gbps, the time available for per-packet decision-making is constrained to less than 100 nanoseconds per packet, which poses challenges for complex computations required for DL-based solutions. In fact, an expensive step in building an ML pipeline is parsing packets for high-speed feature extraction (Unsenix2023-IDP-programmable-switch-flow).
@@ 3.4 Many models; convergence on the horizon
Network traffic modeling began more than two decades ago, resulting in the development of hundreds of models for NetS&P tasks, thanks to the continuous advancement in the field of Machine Learning. The improved capabilities of the ML algorithms, specifically DL models, have allowed researchers in the NetS&P community to train with bigger datasets and larger feature space, while also defining features for various tasks.
Given that network protocols are clearly defined, and substantial progress has been made over the past decades in developing ML-based solutions for NetS&P tasks, it can be argued that the fundamental features useful for modeling are well understood. For illustration, consider a set of research works across the aforementioned NetS&P tasks—anomaly detection (KitSUne-2018-NDSS; GEE-2019), flow classification for botnet detection (FlowLens-NDSS-2021), muti-class attack classification on IoT devices (ADEPT-2021), multi-tab website fingerprinting (multi-tab-FP-transformer-SP2023), flow-correlation attack on Tor networks (oh2022deepcoffea), video fingerprinting (schuster2017beauty; sabzi2024netshaper), dApp fingerprinting (dapp-GNN-FP-2021), prompt inference (prompt-infer-USENIX-SEC-2024) and Tor de-anonymization attack (RECTor-2025)—the raw or fundamental packet-level features at the finest granularity used are packet-size, inter-arrival time (IAT) between packets, direction, protocol, and representations for src/dst IP addresses and port numbers. Features at a coarser granularity, such as flow-level features, are derived from these raw features; for example, flow volume (duration) is the sum of sizes (IAT) of packets in that flow [^1] To be sure, there are a few exceptions. For example, DGA detection is based on DNS payloads, which however would not be available when encrypted DNS protocols are employed. Another exception is TLS feature extraction for representing session-level information. However, TLS is ubiquitous, and the transition from TLS 1.2 to TLS 1.3 has made most features unavailable, thanks to the encryption of handshakes and introduction of ECH. The estimation of certificate size that might be useful for C&C detection (C2-TLS-RAID-2024), could be learned by training on packet-sequences with size being one of the features..

### v4

# AI-based Traffic Modeling for Network Security and Privacy: Challenges Ahead

@@ Abstract
Network traffic analysis using AI (machine learning and deep learning) models made significant progress over the past decades. Traffic analysis addresses various challenging problems in network security, ranging from detection of anomalies and attacks to countering of Internet censorship. AI models are also developed to expose user privacy risks as demonstrated by the research works on fingerprinting of user-visiting websites, IoT devices, and different applications, even when payloads are encrypted.
Despite these advancements, significant challenges remain in the domain of network traffic analysis to effectively secure our networks from evolving threats and attacks. After briefly reviewing the relevant tasks and recent AI models for traffic analysis, we discuss the challenges that lie ahead.
@@ 1 Introduction
The sophistication of networks attacks has increased over time due to the rapid proliferation of new technologies, applications, network protocols and devices. Meanwhile, network traffic keeps increasing, both in volume and speed, further complicating security challenges. To effectively secure the networks, traffic analysis has long been recognized as a critical component. As payloads in network packets are hardly available for signature-based detection with the high adoption of TLS by different applications (Google-transparency-mail; Google-transparency-HTTP), statistical models are becoming inevitable to analyze network traffic. Decades of research efforts have led to the development of various statistical, machine learning (ML) and deep learning (DL) models of different capabilities to tackle multiple network security and privacy (NetS&P) tasks.
DL models, in particular, are valuable for network traffic analysis in NetS&P tasks for two key reasons, or rather, to address two significant challenges. First, network data is enormous; for example, a 1 Gbps link generates hundreds of gigabytes of data per day under high utilization, even when considering only packet headers (excluding the payload). And network bandwidth has increased steadily, with even consumer (broadband) bandwidth now reaching 1 Gbps. Meanwhile, enterprises and telcos/ISPs utilize bandwidth capacities that are orders of magnitude higher. Second, the number of features that can be extracted from network traces can potentially reach several hundreds or even thousands in count. For instance, when modeling a user browsing session as a data point for training/inference, a small session can easily consist of 200 packets, with each packet containing 10 attributes (e.g., packet size, inter-arrival time, and so on). Consequently, representing this relatively short session requires 2,000 features. As we know, DL models are able to process and learn useful patterns from huge datasets with large feature space.
In light of the current state-of-the-art of AI in network traffic modeling, this work discusses the key challenges that lie ahead in securing networks and protecting user privacy.
@@ 2 NetS&P: Tasks and Models
We discuss important NetS&P tasks and AI-based solutions in the literature, without aiming for exhaustiveness.
@@ 2.1 Anomaly detection
AI models: The primary challenge for this task stems from the diverse types of network anomalies, including those that are previously unknown. For instance, while the recent encrypted DNS protocols like DoH provide better privacy for users, they also offer malware developers new ways to hide their communications (survey-encrypted-DNS-mw-2022), say, for exfiltrating sensitive information (real-time-DNS-exfiltration-NDSS-2024). These are relatively new anomalies unknown before DoH was deployed. Existing research in literature tackles this challenge of anomaly detection by developing unsupervised and semi-supervised models. Early research works explored PCA (Principal Component Analysis), statistical hypothesis testing and regression models to detect anomalous patterns in network traffic (PCA-Diot-2004; brauckhoff2009applying; NADA-2018; evidence-gathering-2017). With the emergence of generative deep learning models for unsupervised problem settings (AD-robust-AE-2017; adversarially-learned-AD-ICDM-2018), we witnessed the development of reconstruction-based unsupervised network anomaly detection solutions in the last decade (KitSUne-2018-NDSS; GEE-2019). For example, GEE (GEE-2019) trains a variational autoencoder (VAE) on noisy benign network traffic sessions to learn its corresponding representation in the latent space, which it subsequently uses to detect deviating behaviors during the inference stage.
@@ 2.2 Attack classification
AI models: While the problem might appear to be a multi-class classification problem, the data available for different attacks vary. DDoS requires simple features that capture the traffic rate, whereas detecting the presence of malware that uses DGA (domain generation algorithm) requires analysis of (plain-text) DNS payload. Therefore, solutions in this space typically train binary classifiers to detect specific attacks. For example, several ML algorithms for DDoS detection are evaluated considering both accuracy and inference latency in (DDoS-eval-latency-ICNP-2024). Similarly, prior works have proposed methods for classifying DGA domains (DGA-Rossow-2024), identifying bot traffic to servers using GAN-based data augmentation (botnet-S&P-2020), and detecting encrypted malware communications over TLS (anderson2017machine) as well as Tor (Tor-malware-CCS-2022).
@@ 2.3 IoT device identification
AI models: IoT device identification has received considerable attention over the past decade, resulting in the proposal of various supervised and semi-supervised approaches for classifying known and unknown devices, respectively. They include several works that are based on conventional ML as well as DL models (2017-profileiot; 2019-DIoT; DEFT-2019; IoT-biLSTM-2020; WU-ZEST). A recent work (usenix-2020-youarewhatyou) demonstrates that devices connected to a WiFi network can be identified by training a DL model on the broadcast/multicast traffic observed.
@@ 2.4 Website fingerprinting (WFP) attack
AI models: Given that a browsing session consists of hundreds or even thousands of packets, this topic has witnessed significant development of deep learning-based attack strategies. In other words, the ability of deep learning models to learn from a large feature space represents a substantial advantage for this task. The literature has interesting WFP attack proposals that use various DL architectures ranging from CNN to AE to sequence models such as LSTM and transformer (NDSS-2018-wfp-lstm; pfp-PETS-2019; varCNN-WFP-2019; Usenix-2022-onlinewebsitefingerprinting; webpage_quic; multi-tab-FP-transformer-SP2023; transformer-WFP-CCS-2023; MIL-WFP-USENIX-2024; Levi-2024). There are also strategies proposed for countering them (gong2020FRONTGLUE; smith2022QCSD; GAN-WFP-defense-SP-2022; siby2023PET; Usenix-2023-website-fingerprinting). These models are evaluated under two scenarios: i) a closed-world (lab) setting, where a classification model is trained to classify $n$ specific monitored websites (each with samples collected over time) and is tested for accurate classification of the same $n$ websites, where $n$ is typically in a few hundreds; ii) a more realistic open-world setting, where the challenge is to identify $n$ monitored websites under the assumption that the user visits $m~(m\gg n)$ unmonitored websites, where $m$ is in tens of thousands. Since tunneling of traffic hides packet headers, (different from the previous tasks) the features used here are the basic (or raw) ones such as packet-size, inter-arrival time (IAT) (or time lag) of packet and direction of packet (as browsing generates bidirectional traffic).
@@ 2.5 Other tasks and models
We also highlight the application of adversarial ML, which aims to generate adversarial samples for evading systems utilizing AI models. In the context of NetS&P, an adversarial ML approach can be exploited not only to bypass AI-based security systems but also to counter threats related to privacy and censorship. For instance, authors in (shenoi2023ipet) adapt generative adversarial perturbations (poursaeed2018generative) for evading IoT fingerprinting solutions that employ AI models. And in the space of website fingerprinting, a defense based on GAN, specifically, WGAN-div (wu2018wasserstein), to evade WFP attacker (discriminator) is studied in (GAN-WFP-defense-SP-2022).
@@ 3.1 Data challenges
One of the most important questions is, do the current AI-based solutions generalize beyond the experiments conducted in controlled lab environments? Addressing this question is challenging due to the lack of high-quality labeled data. Most research on network anomaly detection, attack classification, and device identification relies on openly available labeled datasets. However, since these datasets are generated in lab environments, they often contain artifacts or ‘bad design smells’ (bad-smells-EuroSP-2024) that raise concerns on model over-fitting and biases. In  (bad-smells-EuroSP-2024), the authors analyze seven highly-cited datasets and provide evidence of such artifacts in network intrusion detection research. For instance, in one dataset with eight attack classes, two basic features exhibit minimal overlap between benign and attack traffic. As a result, a simple perturbation of these two features—unrelated to the inherent characteristics of attacks—can evade multiple ML and DL models. This simple approach outperforms a state-of-the-art adversarial attack (adversarial-NIDS-2022) employing evolutionary computation and generative adversarial networks. This finding suggests that evaluations in (adversarial-NIDS-2022) may not provide a “meaningful measure of the attack’s effectiveness” (bad-smells-EuroSP-2024) due to dataset flaws. Similar biases in network datasets have also been highlighted for other NetS&P tasks, e.g., website fingerprinting (data-explainable-WFP-PETS-2023; retrace-Tor-2024; genuineTorTraces-2024).
The issue at hand is not about the existing proposals per se, but rather the quality of the datasets used for evaluating network security tasks. Going one step further, the fundamental challenge is the difficulty of validating the quality network traffic. The traffic generated by a single day’s communications in an enterprise network can reach hundreds of gigabytes, and when considering this scale over multiple days, across various enterprises and different verticals, the volume becomes staggering. Therefore, collecting and labeling real-life network traffic flows is an obstacle. Furthermore, real-world network traffic contains sensitive user information, and sharing this data poses significant privacy risks and ethical concerns. Model-based attacks can extract hidden patterns as fingerprinting attacks in the previous section reveal; meta-data from packet headers (excluding payloads) is sufficient to infer sensitive information. Consequently, it is understandable that enterprises and telco operators refrain from publicly sharing traffic collected from their networks for research and development purposes.
@@ 3.2 Practical deployment
Real-time per-packet inference in network traffic analysis faces several practical limitations. Enterprise networks operate at tens of Gbps whereas telco networks operate at 100s of Gpbs. Even at 10 Gbps, the time available for per-packet decision-making is constrained to less than 100 nanoseconds per packet, which poses challenges for complex computations required for DL-based solutions. In fact, an expensive step in building an ML/DL pipeline is parsing packets for high-speed feature extraction (Unsenix2023-IDP-programmable-switch-flow).
@@ 3.4 Many models; convergence on the horizon
For more than two decades, network traffic modeling progressed alongside the advancements in AI, resulting in the development of hundreds of models for NetS&P tasks. We moved from simple statistical methods to ML classifiers such as Decision Trees to more complex and powerful DL models such as transformers; this evolution has been instrumental in training models with bigger datasets and larger feature space, allowing researchers to also define numerous features for various tasks.
Given that network protocols are clearly defined, and substantial progress has been made over the past decades in developing AI-based solutions for NetS&P tasks, it can be argued that the fundamental features useful for modeling are well understood. For illustration, consider a set of research works across the aforementioned NetS&P tasks—anomaly detection (KitSUne-2018-NDSS; GEE-2019), flow classification for botnet detection (FlowLens-NDSS-2021), muti-class attack classification on IoT devices (ADEPT-2021), multi-tab website fingerprinting (multi-tab-FP-transformer-SP2023), flow-correlation attack on Tor networks (oh2022deepcoffea), video fingerprinting (schuster2017beauty; sabzi2024netshaper), dApp fingerprinting (dapp-GNN-FP-2021), prompt inference (prompt-infer-USENIX-SEC-2024) and Tor de-anonymization attack (RECTor-2025)—the raw or fundamental packet-level features at the finest granularity used are packet-size, inter-arrival time (IAT) between packets, direction, protocol, and representations for src/dst IP addresses and port numbers. Features at a coarser granularity, such as flow-level features, are derived from these raw features; for example, flow volume (duration) is the sum of sizes (IAT) of packets in that flow [^1] To be sure, there are a few exceptions. For example, DGA detection is based on DNS payloads, which however would not be available when encrypted DNS protocols are employed. Another exception is TLS feature extraction for representing session-level information. However, TLS is ubiquitous, and the transition from TLS 1.2 to TLS 1.3 has made most features unavailable, thanks to the encryption of handshakes and introduction of ECH. The estimation of certificate size that might be useful for C&C detection (C2-TLS-RAID-2024), could be learned by training on packet-sequences with size being one of the features..